In a vacuum, two particles have charges of q_1 and
Answer #1

F = k × q₁ × q₂ / r²
where
F = force between the particles = 3.3 N
k = constant of coulomb = 9×10^9 Nm²/C²
q₁ = charge on particle 1 = 2.9×10^-6 C
q₂ = charge on particle 2 = ?
r = distance between the particles = 0.31 m
so
3.3 = 9×10^9 ×2.9×10^-6 × q₂ / 0.31²
q₂ = (3.3 × 0.31²) / (9×10^9 × 2.9×10^-6)
q₂ = 1.21×10^-5 C

It is an ATTRACTIVE force, so the charges must have DIFFERENT SIGNS.
q₁ is positive, so q₂ must be negative.

q₂ = - 1.21×10^-5 C